# A Method for On‑Orbit Calibration of the VLAST‑P Electromagnetic Calorimeter

### v1

@@ Abstract
Very Large Area Gamma-ray Space Telescope Pathfinder (VLAST-P), as the technology validation satellite for VLAST mission, is designed to observe high-energy solar bursts on orbit. The CsI electromagnetic calorimeter (ECAL) is one of the key sub-detectors of VLAST-P. In order to study the on-orbit energy calibration method of ECAL, the simulation of VLAST-P was carried out based on Geant4. The results indicated an energy resolution better than 10 % in the 0.1–5 GeV range and a linearity deviation below 2 %, while a dedicated minimum-ionization-particle (MIP) calibration method was developed both to ensure accurate energy reconstruction and to monitor detector stability throughout the in-orbit calibration period.
@@ 1 Space-Based High-Energy Gamma-Ray Observatories
High-energy gamma-ray astronomy has evolved through two main approaches: space-based satellites that directly detect gamma rays and ground-based instruments that observe the particle showers these rays produce in the atmosphere. Ground-based techniques are generally more effective for very high-energy (VHE) gamma rays, as they are less constrained by size and cost, whereas space-based satellites are better suited for observations in the X-ray and high-energy (HE, 30 MeV-30 GeV) ranges [1].
@@
The Compton Gamma Ray Observatory (CGRO), one of the NASA’s gamma ray observatories in the 1990s, was equipped with four gamma-ray telescopes that collectively covered a broad energy range from the keV to GeV scale. For gamma rays with energies above 10 MeV, where pair production becomes the dominant interaction mechanism, CGRO’s high-energy observations were primarily performed by the Energetic Gamma Ray Experiment Telescope (EGRET) [2]. EGRET was designed as a high-energy gamma-ray detector featuring a multilayer spark chamber system for tracking, triggered by a scintillator coincidence system, which was sensitive in the energy range 20 MeV to 30 GeV. Energy measurement was carried out using a total absorption shower counter based on NaI(Tl) crystal, with an overall thickness of 8 radiation length. EGRET achieved an energy resolution of 20–25% across most of its sensitive energy range [3].
@@
The Astrorivelatore Gamma ad Immagini LEggero (AGILE) was developed by the Italian Space Agency (ASI) for high-energy astrophysics research. The AGILE payload included four active detectors that together covered a wide energy range, from hard X-rays to gamma rays. Among them, the gamma-ray imaging detector was made up of a segmented anti-coincidence system, a silicon tracker, and a non-imaging CsI(Tl) mini-calorimeter [4].
@@
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope is a pair-conversion gamma-ray detector with a wide field of view (2.4 sr) and an energy range from below 20 MeV to above 300 GeV. It consists of a silicon-strip tracker, a hodoscopic CsI(Tl) calorimeter, and a segmented anti-coincidence detector. The LAT is designed to measure photon direction, energy, and arrival time while rejecting cosmic-ray background. Since its launch, it has enabled all-sky surveys, source localization, spectral measurements, and studies of gamma-ray bursts, diffuse backgrounds, and potential dark matter signals.
@@
In the near future, The Very Large Area gamma-ray Space Telescope (VLAST) [5, 6] is a proposed mission designed to observe gamma rays across a broad energy spectrum, from MeV to TeV. It will utilize two primary detection methods: Compton scattering and electron-positron pair production. The telescope will operate in low Earth orbit, conducting a detailed survey of gamma-ray sources. Its instrumentation includes: An anti-coincidence detector (ACD) to filter out background events, A tracker detector that also functions as a low-energy calorimeter and a high energy imaging calorimeter (HEIC) for precise measurements.
@@
The VLAST detector system builds on previous established gamma-ray detector designs while making important upgrades. First, the ACD blocks unwanted charged particle events and reduces interference from high-energy events. The silicon tracker and low energy gamma-ray detector (STED) uses thin CsI tiles instead of traditional tungsten plates. This change lets the telescope detect both Compton scattering and pair production events while accurately tracking gamma-ray paths. Finally, the HEIC does three main jobs: it measures particle energies, studies particle showers in detail, and helps distinguish electron events from proton events while estimating their directions.
@@ 2 Very Large Area Gamma-ray Space Telescope Pathfinder
The VLAST-Pathfinder (VLAST-P), a compact version of the VLAST detector, is scheduled for launch in 2026. As a pathfinder mission, it serves dual purposes: validating key technologies for the full VLAST mission, and conducting groundbreaking solar observations. The primary scientific objectives of the VLAST-Pathfinder mission are to address critical gaps in solar high-energy observations by detecting gamma rays in the 100 MeV-GeV energy range, capturing high-energy proton signatures during solar flares, and conducting coordinated observations with China’s ASO-S ("Kuafu-1") solar satellite to enable comprehensive multi-wavelength studies. A key focus will be investigating the fundamental relationship between electron and proton acceleration mechanisms during flare events, which remains poorly understood in current solar physics research.
@@
Figure 1: The schematic of the VLAST-P detector system. It inherits the baseline design of the VLAST detector system. ACD works to distinguish charged particles from gamma rays. Tracker is used to measure the momentum of the charged particles either from the deep universe or from the conversion electron pairs. ECAL is used to measure energy of particles falling into the crystal regions.
@@
Figure 1 shows the schematics of the whole detector system. It basically maintains the design of VLAST detector system. Its detector system comprises three main sub-detector systems: ACD, Si-Tracker and electro-magnetic calorimeter (ECAL). The ACD of VLAST-P is designed with two primary technical objectives: achieve effective separation between charged particles and gamma-ray (detection efficiency better than 99.9%), work stably at high event rates (up to $10\,\mathrm{kHz/m^{2}}$). Tracker detector is built with two main components: CsI convertor and Si-tracker plates. The CsI converter is placed upon Si-tracker plates after ACD. As the name suggests, it aims at converting the incoming photons to a pair of electron and position through photon conversion process. The benefit of utilizing a sensitive detector as convertor instead of a traditional tungsten convertor is that one could collect the energy deposited in CsI which completes the energy measurements. It could be of vital importance in low energy particle reconstructions.
@@
The main role of the detector for particle energy measurement is the electromagnetic calorimeter (ECAL). The ECAL design is limited by the area and weight available on the satellite. It is built with 25 CsI crystal bars placed evenly in a $5\times 5$ array, as shown in the lower part of figure 1. Thallium-doped cesium iodide is chosen as the sensitive material because of its high light yield and good uniformity. The size of each CsI(Tl) crystal is $60\,\mathrm{mm}\times 60\,\mathrm{mm}\times 200\,\mathrm{mm}$; therefore, the total size of the VLAST-P ECAL is approximately $300\,\mathrm{mm}\times 300\,\mathrm{mm}\times 200\,\mathrm{mm}$. The depth and area are optimized to fully exploit the detection capability within the constraints of weight and space.
@@ 3.1 Geant4 Simulation Framework
In this work, the Geant4 toolkit [7, 8] is employed to perform detailed Monte Carlo simulations of particle interactions within the VLAST-P payload. Geant4 provides a comprehensive framework for modeling the transport and interaction of high-energy particles with detector materials, incorporating both electromagnetic and hadronic processes across a broad energy range. Its modular architecture allows for the accurate implementation of geometry, material composition, and detector response, ensuring a realistic description of the experimental setup. For the present study, the simulation focuses on the response of the CsI(Tl)-based electromagnetic calorimeter to cosmic-ray protons, helium nuclei, and solar $\gamma$-rays, which constitute the dominant sources of background and calibration signals in orbit. The use of Geant4 thus enables quantitative evaluation of detector performance, including energy deposition, shower development, and calibration stability, providing critical input for both event reconstruction and on-orbit calibration strategies.
3.2 Detector Geometry and Materials
Figure 2 shows the explosive display of our simulated detector system with front-end tool JSROOT [9]. As shown in the picture, two ACD plates are adopted in the front part of our detector in order to improve the efficiency of telling charged particles from neutral particles like photons. Four CsI(Tl) plates are used to build the converter detector before the Si-tracker plates. In the design of Si-trackers, 6 silicon microstrip plates are used to build the tracking system with 3 heading x and 3 heading y. Each of the two adjacent plates constitutes a layer that provides position information for the hit.
@@ 3.2 Detector Geometry and Materials
As mentioned before, CsI(Tl) crystals are used in the structure of ECAL. It comprises a $5\times 5$ array of crystal arrays. Each crystal is wrapped with a titanium dioxide ($\mathrm{TiO_{2}}$) reflective coating, which serves two purposes: preventing hygroscopic degradation of the crystal and reflecting internally generated scintillation light toward the avalanche photodiode (APD) on one end for optimal photon collection efficiency. The simulation framework enables detailed modeling of detector materials in the VLAST-P detector system, including its carbon fiber support structure and embedded silicone rubber buffer pads. The carbon fiber framework ensures structural stability during launch and orbital operations, while the silicone rubber layers mitigate micro-vibrations that could probably affect crystal performance.
@@ 3.3 Performance and Validation
In the simulation study of the ECAL, one of the most important metrics is the energy resolution. Gamma rays with energies ranging from $20\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$ were simulated to be incident directly on the detector system. The energy of each event is reconstructed by summing the energies of adjacent crystals surrounding the one with the maximum deposition. Tracker information is used to determine whether the incident photon converted into an electron–positron pair. The energy distribution at each incident energy point is fitted with a Crystal Ball function to extract the mean reconstructed energy and the full width at half maximum (FWHM). The energy resolution $R$ is calculated from the FWHM of the energy peak as follows:
@@
where $\mu$ is the mean energy of the peak. The factor 2.35 converts the FWHM to the standard deviation, assuming a Gaussian distribution.
(a) Energy linearity
(b) Energy resolution
@@
The performance of the VLAST-P ECAL was evaluated in terms of both energy linearity and resolution for incident gamma rays. As shown in figure 3(a), the measured energy deposited in the CsI(Tl) scintillator exhibits a nearly proportional response to the true incident energy over the range from $20\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, demonstrating good energy linearity with less than 10% deviations. Correspondingly, the energy resolution, illustrated in figure 3(b), remains around 5% across a wide energy range when gamma rays are directly incident on the detector. This is due to the CsI(Tl) material absorbing most of the energy of the incoming particles, with only a small fraction of energy leakage occurring at higher energies, ensuring reliable energy reconstruction.

### v2

@@ Abstract
Very Large Area Gamma-ray Space Telescope Pathfinder (VLAST-P), serving as the technology validation satellite for VLAST mission, is designed to observe high-energy solar bursts on orbit. The CsI electromagnetic calorimeter (ECAL) is one of the key sub-detectors of VLAST-P. In order to study the on-orbit energy calibration method of ECAL, the simulation of VLAST-P was carried out based on Geant4. The results indicated an energy resolution better than 10 % in the 0.1–5 GeV range, while a dedicated minimum-ionization-particle (MIP) calibration method was developed both to ensure accurate energy reconstruction and to monitor detector stability throughout the in-orbit calibration period.
@@ 1 Space-Based High-Energy Gamma-Ray Observatories
High-energy gamma-ray astronomy has evolved through two main approaches: space-based satellites that directly detect gamma rays and ground-based instruments that observe the particle showers these rays produce in the atmosphere. Ground-based techniques are generally more effective for very high-energy (VHE) gamma rays, as they are less constrained by size and cost, whereas space-based satellites are better suited for observations in the X-ray and high-energy (HE, 30 MeV-30 GeV) ranges [Gamma-ray_Astronomy].
@@
The Compton Gamma Ray Observatory (CGRO), one of NASA’s gamma ray observatories in the 1990s, was equipped with four gamma-ray telescopes that collectively covered a broad energy range from the keV to GeV scale. For gamma rays with energies above 10 MeV, where pair production becomes the dominant interaction mechanism, CGRO’s high-energy observations were primarily performed by the Energetic Gamma Ray Experiment Telescope (EGRET) [Thompson2008]. EGRET was designed as a high-energy gamma-ray detector featuring a multilayer spark chamber system for tracking, triggered by a scintillator coincidence system, which was sensitive in the energy range 20 MeV to 30 GeV. Energy measurement was carried out using a total absorption shower counter based on NaI(Tl) crystal, with an overall thickness of 8 radiation length. EGRET achieved an energy resolution of 20–25% across most of its sensitive energy range [EGRETCalib].
@@
The Astrorivelatore Gamma ad Immagini LEggero (AGILE) was developed by the Italian Space Agency (ASI) for high-energy astrophysics research. The AGILE payload included four active detectors that together covered a wide energy range, from hard X-rays to gamma rays. Among them, the gamma-ray imaging detector was made up of a segmented anti-coincidence system, a silicon tracker, and a non-imaging CsI(Tl) mini-calorimeter [Vercellone2024pgh].
@@
The Large Area Telescope (LAT) onboard the Fermi Gamma-ray Space Telescope is a pair-conversion gamma-ray detector with a wide field of view (2.4 sr) and an energy range from below 20 MeV to more than 300 GeV [FermiLAT]. It consists of a silicon-strip tracker, a hodoscopic CsI(Tl) calorimeter, and a segmented anti-coincidence detector. The LAT is designed to measure photon direction, energy, and arrival time while rejecting cosmic-ray background. Since its launch, it has enabled all-sky surveys, source localization, spectral measurements, and studies of gamma-ray bursts, diffuse backgrounds, and potential dark matter signals.
@@
In the near future, The Very Large Area gamma-ray Space Telescope (VLAST) [2022AcASn..63...27F, Pan_2024] is a proposed mission designed to observe gamma rays across a broad energy spectrum, from MeV to TeV. It will utilize two primary detection methods: Compton scattering and electron-positron pair production. The telescope will operate in low Earth orbit, conducting a detailed survey of gamma-ray sources. Its instrumentation includes: an anti-coincidence detector (ACD) to filter out background events, a tracker detector that also functions as a low-energy calorimeter and a high energy imaging calorimeter (HEIC) for precise measurements.
@@
The VLAST detector system builds on previously established gamma-ray detector designs while making important upgrades. First, the ACD blocks unwanted charged particle events and reduces interference from high-energy events. The silicon tracker and low energy gamma-ray detector (STED) uses thin CsI tiles instead of traditional tungsten plates. This change lets the telescope detect both Compton scattering and pair production events while accurately tracking gamma-ray paths. Finally, the HEIC performs three main functions: it measures particle energies, studies particle showers in detail, and helps distinguish electron events from proton events while estimating their directions.
@@ 2 Very Large Area Gamma-ray Space Telescope Pathfinder
The VLAST-Pathfinder (VLAST-P), a compact version of the VLAST detector, is scheduled for launch in 2026. As a pathfinder mission, it serves dual purposes: validating key technologies for the full VLAST mission, and conducting groundbreaking solar observations. The primary scientific objectives of the VLAST-Pathfinder mission are to address critical gaps in solar high-energy observations by detecting gamma rays in the 100 MeV-5 GeV energy range, capturing high-energy proton signatures during solar flares, and conducting coordinated observations with China’s ASO-S ("Kuafu-1") solar satellite to enable comprehensive multi-wavelength studies. A key focus will be on investigating the fundamental relationship between electron and proton acceleration mechanisms during flare events, which remains poorly understood in current solar physics research.
@@
Figure 1: The schematics of the VLAST-P detector system. It inherits the baseline design of the VLAST detector system. ACD serves to distinguish charged particles from gamma rays. The tracker is used to measure the tracks of charged cosmic rays or converted electron pairs. The ECAL is used to measure the energy of particles entering the crystals.
@@
Figure 1 shows the schematics of the whole detector system. It maintains the design of VLAST detector system. Its detector system comprises three main sub-detector systems: ACD, tracker and electro-magnetic calorimeter (ECAL). The ACD of VLAST-P is designed with two primary technical objectives: to achieve effective separation between charged particles and gamma rays (with a charged particle detection efficiency exceeding 99.9%), maintain optimal performance at high event rates (up to $10\,\mathrm{kHz/m^{2}}$). Tracker detector is built from two main components: a 15 mm thick CsI converter and a silicon tracker. The CsI converter is placed upon silicon tracker after ACD. It acts as both a photon converter that converts the incoming photons to a pair of electron and positron and an active calorimeter with radiation length of around $0.8\,X_{0}$. The benefit of utilizing a sensitive detector as converter instead of a traditional tungsten converter is that one could collect the energy deposited in CsI which completes the energy measurements. It could be of vital importance in low energy particle reconstructions. Following the converter, the silicon tracker consists of six silicon micro-strip layers arranged in an alternating x-y configuration. Each pair of orthogonal strips provides 2D coordinate measurements. This design characterizes the electron-positron pairs from photon conversion with a spatial resolution better than $50\,\mathrm{um}$ [AZZARELLO2016378].
@@
The main sub-detector for particle energy measurement is the electromagnetic calorimeter (ECAL). The detector is optimized to cover an energy range of $100\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, although the ECAL design is limited by the area and weight available on the satellite. It is built with 25 CsI(Tl) crystal bars arranged in a $5\times 5$ array, as shown in the lower part of figure 1. Thallium-doped cesium iodide is chosen as the sensitive material because of its high light yield and good uniformity.
The size of each CsI(Tl) crystal is $60\,\mathrm{mm}\times 60\,\mathrm{mm}\times 200\,\mathrm{mm}$. With a radiation length ($X_{0}$) of $1.86\,\mathrm{cm}$ for CsI(Tl), the longitudinal depth of the ECAL corresponds to approximately $10.8\,X_{0}$, which is sufficient to contain a large portion of the electromagnetic showers from GeV photons. The crystals are wrapped with $\mathrm{TiO_{2}}$ reflective films to fully collect the scintillation photons. The crystals are separated by a mechanical gap of $1\,\mathrm{mm}$; thus, the total active volume is approximately $300\,\mathrm{mm}\times 300\,\mathrm{mm}\times 200\,\mathrm{mm}$.
The required readout energy range is designed to cover the energy range up to $5\,\mathrm{GeV}$. To measure incident photons up to $5\,\mathrm{GeV}$, simulations indicate a maximum energy deposition of approximately $4.1\,\mathrm{GeV}$ in a single crystal, which determines the upper limit of the measurement range. At the low-energy end, the energy threshold of single crystal in simulation is set to $3\,\mathrm{MeV}$; therefore, the electronic noise must be controlled to meet the threshold. As a result, the readout system is required to cover a wide dynamic range from the $3\,\mathrm{MeV}$ up to the $4.1\,\mathrm{GeV}$ saturation level.
@@ 3.1 Geant4 Simulation Framework
In this work, the Geant4 toolkit [ALLISON2016186, AGOSTINELLI2003250] is employed to perform detailed Monte Carlo simulations of particle interactions within the VLAST-P payload. Geant4 is a comprehensive framework for modeling the transport and interaction of high-energy particles with detector materials, incorporating both electromagnetic and hadronic processes across a broad energy range. Its modular architecture allows for the accurate implementation of geometry, material composition, and detector response, ensuring a realistic description of the experimental setup. For the present study, the simulation focuses on the response of the CsI(Tl)-based electromagnetic calorimeter to cosmic-ray protons, helium nuclei, and solar $\gamma$-rays, which constitute the dominant sources of background and calibration signals in orbit. The use of Geant4 thus enables quantitative evaluation of detector performance, including energy deposition, shower development, and calibration stability, providing critical input for both event reconstruction and on-orbit calibration strategies.
3.2 Detector Simulation Setup
Figure 2 shows the exploded view of our simulated detector system drawn with front-end tool JSROOT [JSROOT]. The components are displayed in separate layers to illustrate the internal structure. As shown in the picture from left to right, the two outermost plates (shown in light green) are the ACD layers, which are positioned at the front to effectively distinguish charged particles from photons. Next to the ACD, four CsI(Tl) plates (shown in yellow) are placed before the six silicon tracker layers (depicted in blue and red layers). Specifically, four CsI(Tl) plates serve as the converter while the six silicon microstrip plates (used to measure x and y coordinates) provide position information with 3 oriented along the x-axis and 3 along the y-axis. Each of the two adjacent plates constitutes a layer that provides position information for the hit. Finally, on the far right, the ECAL is visualized as the $5\times 5$ array of purple crystal bars, which are used to measure the energy of particles entering the crystals. Detailed mechanical and electronic components are also visible: the small red squares attached to the crystal ends are APDs, while the purple frame-like structures correspond to the silicone rubber pads designed for shock absorption. Inclusion of these fine details ensures a more realistic representation of the detector geometry.
@@ 3.2 Detector Simulation Setup
As mentioned before, CsI(Tl) crystals are used in the structure of ECAL. It comprises a $5\times 5$ array of crystal arrays. Each crystal is wrapped with a titanium dioxide ($\mathrm{TiO_{2}}$) reflective coating, which serves two purposes: preventing hygroscopic degradation of the crystal and reflecting internally generated scintillation light toward the avalanche photodiode (APD) on one end for optimal photon collection efficiency. The simulation framework enables detailed modeling of detector materials in the VLAST-P detector system, including its carbon fiber support structure and embedded silicone rubber buffer pads. The carbon fiber frame ensures structural stability during launch and orbital operations, while the silicone rubber layers mitigate micro-vibrations that could probably affect crystal performance.
In addition to the geometry and materials, the simulation incorporates a digitization process to process the raw energy deposition to obtain a more realistic reconstructed energy value. The process begins by converting the deposited energy into the number of photoelectrons based on the light yield of the CsI(Tl) crystals. Subsequently, Gaussian smearing is applied to the signal to account for both the statistical fluctuations and the non-uniformity of the crystal response, where the latter was determined to be $2\%$ based on our single-crystal cosmic ray tests. The total signal also includes the ionization energy deposited by particles passing directly through the APDs. Finally, an electronic noise level of approximately $1\,\mathrm{MeV}$ is superimposed on the signal to mimic the realistic detector performance.
@@ 3.3 Performance and Validation
By employing a full Geant4-based simulation, the intrinsic energy response of the calorimeter to monoenergetic photons was evaluated. Gamma rays with energies ranging from 50 MeV to 5 GeV were simulated at normal incidence to the detector surface. The primary particles were emitted from a planar source of size $150\penalty 10000\ \mathrm{mm}\times 150\penalty 10000\ \mathrm{mm}$, positioned 10 mm away from the center of the ACD’s front surface. The particles were incident perpendicularly to the detector surface. A total of $1\times 10^{5}$ events were generated to ensure sufficient statistical precision for each incident energy point. The energy of each event is reconstructed by summing the energies of adjacent crystals surrounding the one with the maximum deposition. The VLAST-P tracker consists of two components: an active converter layer, where incident gamma rays are expected to undergo pair production, and a silicon tracking section that records the subsequent electron-positron trajectories with high spatial resolution. The tracker serves two essential functions: (i) photon conversion, by providing an active medium in which gamma rays can convert into an electron–positron pair, and (ii) direction reconstruction, by accurately measuring the pair trajectories to infer the incoming photon’s direction. In the context of the ECAL simulation, tracker information is used to identify whether the photon has converted in the tracker; events that convert upstream are either rejected or treated separately, ensuring that the energy response to calorimeter is evaluated only for photons entering the ECAL as intended. As illustrated in figure 3(a), the energy distribution at each incident energy point is fitted with a Crystal Ball function [PhysRevD.34.711] to extract the mean reconstructed energy and the full width at half maximum (FWHM). The energy resolution $R$ is calculated from the FWHM of the energy peak as follows:
@@
where $\mu$ is the fitted energy of the peak. The factor 2.35 converts the FWHM to the standard deviation, assuming a Gaussian distribution.
(a) Example fit at 1000 MeV
(b) Energy linearity
(c) Energy resolution
@@
The performance of the VLAST-P ECAL was evaluated in terms of both energy linearity and resolution for incident gamma rays. As shown in figure 3(b), the measured energy deposited in the CsI(Tl) scintillator exhibits a nearly proportional response to the true incident energy over the range from $20\,\mathrm{MeV}$ to $5\,\mathrm{GeV}$, demonstrating good energy linearity with less than 10% deviations. Correspondingly, the energy resolution, illustrated in figure 3(c), remains around 5% across a wide energy range for gamma rays normally incident on the detector. This is due to the high stopping power of the CsI(Tl) crystals, which absorb most of the incident energy. Consequently, only a small fraction of energy leakage occurs even at higher energies, ensuring accurate energy reconstruction.